find the coordinates of the x intercept and the y intercept of this linear equation 3x-4y=24
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

find the coordinates of the x intercept and the y intercept of this linear equation 3x-4y=24
----
If x = 0, y = -6
If y = 0, x = 8
--------
Cheers,
Stan H.

Answer by rfer(16322)

3x-4y=24
3x-4*0=24
x=8
(8,0)
3*0-4y=24
y=-6
(0,-6)